# Learning Robust and Correct Controllers Guided by Feasibility-Aware Signal Temporal Logic via BarrierNet

### v2

@@ IV-D Learning Robust Controllers
Figure 1: Overall structure of the controller. Purple modules run only at $t=0$, while blue modules operate at every later time step. The dashed box outlines the controller $\pi(\cdot,\boldsymbol{\theta})$ ((2) or (3)) over one rollout (epoch).
Theorem 5 shows that maintaining a feasibility-aware STL robustness above zero guarantees both (i) satisfaction of the STL specification and (ii) feasibility of the underlying QP at all times. With this guarantee, the BarrierNet architecture in (10), integrated with a model-based RL framework, can be used to synthesize a controller that enforces the feasibility-aware robustness during execution while gradually improving it through training. In this subsection, we first explain why in (10), $b(\mathbf{x},t,\bm{\theta}_{b},\mathbf{x}_{0})$ and $\mathbf{P}_{\text{inip}}(\mathbf{x}_{0},\boldsymbol{\theta}_{\text{inip}})$ all depend on the initial condition $\mathbf{x}_{0}$. Then we describe the structure of the entire neural network controller $\pi(\mathbf{x}(t),\bm{\theta})$ in (2) or $\pi(\mathbf{x}_{0:t},\bm{\theta})$ in (3). Finally, we introduce the training process of the controller.
Hyperparameters Depending on the Initial Condition. Consider a predicate in Category II or III with the associated HOCBF $b(\mathbf{x},t)=h(\mathbf{x})+\gamma(t,\boldsymbol{\omega})$. The constraints (22a) and (24a) that shape the function $\gamma(t,\boldsymbol{\omega})$ explicitly depend on the initial state $\mathbf{x}_{0}$. Therefore, the hyperparameter vector $\boldsymbol{\omega}$ must be adapted to each initial condition rather than fixed globally.
@@
To ensure that the HOCBF constraint hierarchy enforces set invariance, we must also satisfy $\psi_{i}(\mathbf{x}_{0},0)\geq 0$ for all $i=1,\ldots,m-1$. Since $b(\mathbf{x}_{0},0)>0$, (9) implies that each $\psi_{i}(\mathbf{x}_{0},0)$ can be made nonnegative by selecting sufficiently large multipliers $p_{i}$ at $t=0$. Because these lower bounds on $p_{i}$ depend on the initial condition $\mathbf{x}_{0}$, we generate them through a neural network that takes $\mathbf{x}_{0}$ as input and is parameterized by $\boldsymbol{\theta}_{\mathrm{inip}}$, then we have $p_{i}=p_{i}(\mathbf{x}_{0},\boldsymbol{\theta}_{\mathrm{inip}})$. We denote the resulting initial multipliers collectively as $\mathbf{p}_{j}=[p_{1,j},\ldots,p_{m-1,j}]^{\top}$ and $\mathbf{P}_{\text{inip}}=[\mathbf{p}_{1}^{\top},\ldots,\mathbf{p}_{M}^{\top}]\in\mathbb{R}^{(m-1)M}$. $\mathbf{P}_{\mathrm{inip}}$ serves as the initialization of $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\mathrm{inip}})$ at $t=0$.
@@
where $\bm{\Omega}=[\bm{\omega}_{1}^{\top},\ldots,\bm{\omega}_{N}^{\top}]\in\mathbb{R}^{2N},N\leq M$ collects all hyperparameters defining the time-varying terms $\gamma(t,\bm{\omega}_{j})$. The network $\mathfrak{N}$ is parameterized by the trainable variables $\bm{\theta}_{b}$ and $\boldsymbol{\theta}_{\mathrm{inip}}$. All constraints on $\gamma$ in (22), (24), and (27) are equivalently enforced as constraints on the output $\bm{\Omega}$ of InitNet. For hyperparameter constraints of the form $\omega\in[\underline{\omega},\overline{\omega}]$, we enforce them by applying a Sigmoid activation at the final layer of $\mathfrak{N}$. For one-sided constraints such as $\omega\in[\underline{\omega},\infty)$ or $\omega\in(-\infty,\overline{\omega}]$, we use a Softplus activation (possibly combined with sign flipping) to guarantee positivity or negativity as required. With these activations, the generated vector $\bm{\Omega}$ automatically satisfies all constraints in (22), (24), and (27). Similarly, for the multipliers $p_{i}$, $i=1,\ldots,m-1$, the initialization constraints
$p_{i}\;>\;\max\!\Big\{-\dot{\psi}_{i-1}(\mathbf{x}_{0},0,\boldsymbol{\omega},\mathbf{p})\big/\alpha_{i}(\psi_{i-1}(\mathbf{x}_{0},0,\boldsymbol{\omega},\mathbf{p})),\;0\Big\}$ (39)
@@
are enforced by applying Softplus activations to the corresponding output channels of $\mathfrak{N}$, ensuring that each $p_{i}$ satisfies its required lower bound. InitNet is only used at time $t=0$ to provide a set of HOCBFs and the initial multipliers for corresponding class $\kappa$ functions, which are then used to train $\bm{\theta}_{b}$ to obtain $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\mathrm{inip}})$.
@@
RefNet Structure. At each discrete time step, we employ a neural network (with or withour memory), referred to as RefNet, parameterized by $(\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{q})$—to generate the reference terms $\mathbf{F}(\mathbf{x}_{0:t},\boldsymbol{\theta}_{f}),\mathbf{Q}(\mathbf{x}_{0:t},\boldsymbol{\theta}_{q})$ for controller $\pi(\mathbf{x}_{0:t},\bm{\theta})$ in (3) or $\mathbf{F}(\mathbf{x}(t),\boldsymbol{\theta}_{f}),\mathbf{Q}(\mathbf{x}(t),\boldsymbol{\theta}_{q})$ for controller $\pi(\mathbf{x}(t),\bm{\theta})$ in (2)—which serve as reference inputs to the BarrierNet module.
@@
Neural Network Controller Structure. The hyperparameters of the time-varying HOCBFs are obtained from the InitNet $\mathfrak{N}(\mathbf{x}_{0},\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{\mathrm{inip}})$, which produces the vectors $\boldsymbol{\Omega}$ and $\mathbf{P}_{\mathrm{init}}$ that encode all $\gamma$-function hyperparameters and the initial multipliers.
@@
Given the outputs of RefNet and InitNet, the BarrierNet layer then solves the differentiable QP in (10) to produce the control input $\mathbf{u}^{*}(t)$ at each time step, while ensuring that the feasibility-aware STL robustness—including all STL requirements and control bounds—is respected along the entire closed-loop trajectory. By applying $\mathbf{u}^{*}(t)$ sequentially to the system (1), BarrierNet rolls out the full closed-loop trajectory, from which the overall robustness $\rho(\tilde{\varphi},\mathbf{x}_{0:t_{T}},0)$ of the augmented STL specification and its associated cost $J(\mathbf{u}_{0:t_{T}-1})$ are evaluated. This trajectory-level evaluation is then used to update the controller parameters $\boldsymbol{\theta}$ during reinforcement training. Overall, the neural network controller $\pi(\cdot,\boldsymbol{\theta})$, where $\boldsymbol{\theta}=(\boldsymbol{\theta}_{q},\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{\text{inip}},\boldsymbol{\theta}_{p}),$ consists of (i) InitNet for initial hyperparameter generation, (ii) RefNet for predicting $\mathbf{F}$ and $\mathbf{Q}$, and (iii) the BarrierNet dQP layer that ensures all feasibility-aware STL robustness constraints are satisfied. The full architecture is illustrated in Fig. 1.
@@
Training Neural Network Controller. Following the procedure in [liu2023safe], we randomly sample $V$ initial conditions $\mathbf{x}_{0}^{v}$, $v=1,\ldots,V$. For each initial state, we roll out the closed-loop system (1) under the controller $\pi$ up to the horizon $T$, thereby generating $V$ state and control trajectories. For each trajectory, we compute the STL robustness and its associated cost $J$, and use their empirical mean to approximate the expectation. Thus, the optimization problem in (8) can be written as
@@
where the superscript $v$ denotes the $v$-th sample. By substituting the system dynamics into the objective, the problem becomes an unconstrained optimization over the parameters of the controller. Since the QP in (10) is differentiable with respect to its inputs and parameters via the OptNet technique [amos2017optnet], we backpropagate the gradient of the objective in (40) through the QP layer to all components of $\bm{\theta}$. The gradients of the STL robustness are computed analytically and automatically using a modified implementation of STLCG [leung2020backpropagation] adapted to the exponential robustness in [liu2023robust]. We then update the parameters using these gradients. At each gradient step, we resample $V$ initial conditions from the set $\mathcal{X}_{0}$ to improve exploration of the initial-state distribution, and we employ the stochastic optimizer Adam [kingma2014adam] to perform the training.
We summarize our solution to Problem 1 in Algorithm 1.
@@ 2 Set up constraints on $\bm{\omega}$ using (22), (24), (27);
3 Set up constraints on $p_{i,j}$ using (9);
@@ 11   Update $\bm{\theta}$ using Adam optimizer;
Corollary 1 (Feasibility-Aware Correctness of BarrierNet Controllers).
Consider the system (1) and the STL specification $\varphi$. Construct the HOCBFs and their feasibility subformulas $\varphi_{j}^{\mathrm{fea}}$ as in Sec. IV-B, and define the augmented STL specification $\widetilde{\varphi}$ by combining $\varphi$ with all feasibility and control-bound subformulas as in (35). Let a neural network controller be given whose last layer is the differentiable QP (10), which enforces all highest-order HOCBF constraints, deletion rules, and control bounds at every time step. If the closed-loop trajectory generated by this controller satisfies $\rho(\widetilde{\varphi},\mathbf{x},0)\geq 0$, then the controller guarantees that the dQP (10) remains feasible for all $t\geq 0$, and the resulting closed-loop trajectory satisfies the original STL specification $\varphi$.
Proof.
This follows immediately from Thm. 5. ∎
Remark 2.
Since we can only feed discrete data into neural networks, the QP (10) is solved in discrete time. Hence, the inter-sampling effect should also be considered to guarantee the correctness of the controller. A possible solution to address this effect is the sampling-aware CBF framework [liu2025sampling]. We will consider this in future work.
@@ V-A Case Study I: Linear Dynamical System — Double Integrator
where $\mathbf{x}=[x\ y\ v_{x}\ v_{y}]^{\top}$ and $\mathbf{u}=[u_{1}\ u_{2}]^{\top}$, with $x\ y]^{\top}$ denoting the 2D position, $v_{x}\ v_{y}]^{\top}$ the velocity, and $u_{1}\ u_{2}]^{\top}$ the acceleration of the robot. During the initialization of the neural network controller, the control inputs are bounded. However, in the subsequent optimization stage, the inputs are treated as unconstrained, while the STL specification still accounts for the actual input limits in (34), and the cost function in (40) employs an $\ell_{2}$ penalty with a weight of $0.003$ to discourage large accelerations.
@@ V-A1 Neural Network Controller without Memory
Comparison setup. We construct the HOCBFs (19), (20) and train the controller proposed in this paper under two different configurations. (i) In the first configuration, the multipliers $p_{i,j}$ are time-varying, while the feasibility subformula (33) is not incorporated; we refer to this controller as BN-VarP. (ii) In the second configuration, the multipliers $p_{i,j}$ are time-varying and the feasibility subformula (33) is enforced; we refer to this controller as FeasiBN-VarP. The position state [xy][x\ y] serves as the input to the trained multiplier $p_{i,j}([x\ y],\boldsymbol{\theta}p,\mathbf{P}{\text{init}})$.
For comparison, we include our previous work [liu2023learning], where the multipliers $p_{i,j}$ remain constant within each rollout; we call this controller BN-FixedP. Another benchmark is [liu2023safe], in which a neural network controller without BarrierNet—i.e., a standard Fully Connected Neural Network (FCNet)—is trained to satisfy an STL task. This is equivalent to directly using the reference control $\mathbf{F}$. We refer to this controller as FCNet. To ensure a fair comparison, we assume that the system dynamics are known for [liu2023safe], and we use the same objective function, optimizer, and neural network architectures; that is, BN-FixedP and FCNet share the same architecture as $\mathbf{F}(\mathbf{x},\bm{\theta}_{f})$.
The training curves for all learning-based methods are shown in Fig. 3. In addition, we apply the method in [lindemann2018control] (extended to the HOCBF setting) without learning: we construct HOCBFs with fixed hyperparameters and solve the QP (10) with $\mathbf{F}=\mathbf{0}$. The hyperparameters are randomly chosen but satisfy all constraints (22), (24), and (27). We refer to this baseline as HOCBF. The resulting average objective values and robustness over 10 random initial conditions are shown in Fig. 3, where the HOCBF baseline is illustrated by pink dashed lines. Representative trajectories obtained by the five approaches under 10 random initial conditions are shown in Fig. 2(b). The inputs over time for the five methods are shown in Fig. 4. We select the second task, $F_{[2,5]}Reg_{2}$ ($j=2$), and the fourth task, $G_{[0,5]}(\neg Obs_{4})$ ($j=4$), to illustrate the corresponding $p_{i,j}$ values ($i=1,2$) for the five methods in Fig. 5.
@@
Figure 3: Learning curves for BN-FixedP, BN-VarP, FeasiBN-VarP, and FCNet. Dashed lines show the results of directly using HOCBFs. (a) mean robustness during training. (b) mean objective values during training.
@@
Analysis and Discussion. From Fig. 3(a), we observe that when using BarrierNet (BN-FixedP, BN-VarP, and FeasiBN-VarP), all three methods achieve high robustness after sufficient training iterations. Specifically, BN-FixedP reaches a maximum robustness of approximately 0.45 during 500 iterations, while BN-VarP and FeasiBN-VarP reach about 0.6 and 0.7, respectively. Moreover, the robustness obtained by each BarrierNet-based method is higher than that of FCNet at nearly every iteration, demonstrating the correctness of Corollary 1. Occasionally, FCNet achieves higher robustness than BN-FixedP. This occurs because fixed multipliers limit the performance of BarrierNet, especially when the STL specification involves control bounds. In contrast, the use of time-varying multipliers in BN-VarP and FeasiBN-VarP significantly improves robustness, confirming that time-varying multipliers enhance the satisfaction of STL specifications, as claimed in Rem. 1. We further note that BN-VarP produces negative robustness at certain iterations, and further inspection reveals that the controller fails to satisfy the control-bound–related specification at those times. In comparison, FeasiBN-VarP maintains positive robustness throughout training, showing that the feasibility-aware formulation effectively guides the QP toward satisfying all constraints.
@@
In both Fig. 4 and Fig. 5, we plot the input and multiplier profiles from a single representative trajectory chosen from the 10 trajectories. From Fig. 4(b), we observe that after training, both the BarrierNet-based methods and FCNet produce control inputs that satisfy the input bounds. The input profiles of BN-FixedP closely resemble those of the HOCBF baseline after training. This is further confirmed in Fig. 5, where the learned multipliers of BN-FixedP are very similar to those of the HOCBFs. These results indicate that fixed multipliers substantially limit the ability of BarrierNet to improve performance. In contrast, BN-VarP and FeasiBN-VarP adjust the inputs earlier and more aggressively to satisfy the specifications, as shown in Fig. 4(b). This behavior is consistent with Fig. 5, where the time-varying multipliers adapt across time steps according to the control needs. Such adaptability enables these methods to achieve higher robustness.

### v3

@@ IV-D Learning Robust Controllers
Corollary 1 shows that maintaining a feasibility-aware STL robustness above zero indicates both (i) satisfaction of the STL specification and (ii) feasibility of the underlying QP at all times. With this monitoring, the BarrierNet architecture in (10), integrated with a model-based RL framework, can be used to synthesize a controller that accounts for the feasibility-aware robustness during execution while gradually improving it through training. In this subsection, we first explain why in (10), $b(\mathbf{x},t,\bm{\theta}_{b},\mathbf{x}_{0})$ and $\mathbf{P}_{\text{inip}}(\mathbf{x}_{0},\boldsymbol{\theta}_{\text{inip}})$ all depend on the initial condition $\mathbf{x}_{0}$. Then we describe the structure of the entire neural network controller $\pi(\mathbf{x}(t),\bm{\theta})$ in (2) or $\pi(\mathbf{x}_{0:t},\bm{\theta})$ in (3). Finally, we introduce the training process of the controller.
Hyperparameters Depending on the Initial Condition. Consider a predicate in Category II or III with the associated HOCBF $b(\mathbf{x},t,\bm{\theta}_{b})=h(\mathbf{x})+\gamma(t,\boldsymbol{\omega})$. The constraints (22a) and (24a) that shape the function $\gamma(t,\boldsymbol{\omega})$ explicitly depend on the initial state $\mathbf{x}_{0}$. Therefore, the hyperparameter vector $\boldsymbol{\omega}$ must be adapted to each initial condition rather than fixed globally.
@@
To ensure that the HOCBF constraint hierarchy enforces set invariance, we must also satisfy $\psi_{i}(\mathbf{x}_{0},0,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})\geq 0$ for all $i=1,\ldots,m-1$. Since $b(\mathbf{x}_{0},0,\bm{\theta}_{b})>0$, (9) implies that each $\psi_{i}(\mathbf{x}_{0},0,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$ can be made nonnegative by selecting sufficiently large multipliers $p_{i}^{\mathrm{inip}}$ at $t=0$. To ensure the satisfaction of the initial feasibility condition, the parameters $p_{i}^{\mathrm{inip}}$ are selected such that each $\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})=\epsilon_{j}>0$, where $\epsilon_{j}$ is a sufficiently small positive constant. Because these lower bounds on $p_{i}^{\mathrm{inip}}$ depend on the initial condition $\mathbf{x}_{0}$, we generate them through a neural network that takes $\mathbf{x}_{0}$ as input and is parameterized by $\boldsymbol{\theta}_{\mathrm{inip}}$, then we have $p_{i}^{\mathrm{inip}}=p_{i}^{\mathrm{inip}}(\mathbf{x}_{0},\boldsymbol{\theta}_{\mathrm{inip}})$. We denote the resulting initial multipliers collectively as $\mathbf{p}_{j}=[p_{1,j}^{\mathrm{inip}},\ldots,p_{m-1,j}^{\mathrm{inip}}]^{\top}$ and $\mathbf{P}_{\text{inip}}=[\mathbf{p}_{1}^{\top},\ldots,\mathbf{p}_{M}^{\top}]\in\mathbb{R}^{(m-1)M}$. $\mathbf{P}_{\mathrm{inip}}$ serves as the initialization of $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\mathrm{inip}})$ at $t=0$.
@@
where $\bm{\Omega}=[\bm{\omega}_{1}^{\top},\ldots,\bm{\omega}_{N}^{\top}]\in\mathbb{R}^{2N},N\leq M$ collects all hyperparameters defining the time-varying terms $\gamma_{j}(t,\bm{\omega}_{j})$. The network $\mathfrak{N}$ is parameterized by the trainable variables $\bm{\theta}_{b}$ and $\boldsymbol{\theta}_{\mathrm{inip}}$. All constraints on $\gamma_{j}$ in (22), (24), and (27) are equivalently enforced as constraints on the output $\bm{\Omega}$ of InitNet. For hyperparameter constraints of the form $\omega\in[\underline{\omega},\overline{\omega}]$, we enforce them by applying a Sigmoid activation at the final layer of $\mathfrak{N}$. For one-sided constraints such as $\omega\in[\underline{\omega},\infty)$ or $\omega\in(-\infty,\overline{\omega}]$, we use a Softplus activation (possibly combined with sign flipping) to guarantee positivity or negativity as required. With these activations, the generated vector $\bm{\Omega}$ automatically satisfies all constraints in (22), (24), and (27). Similarly, for the multipliers $p_{i}^{\mathrm{inip}}$, $i=1,\ldots,m-1$, the initialization constraints
$p_{i}^{\mathrm{inip}}>\max\!\Big\{-\dot{\psi}_{i-1}(\mathbf{x}_{0},0,\boldsymbol{\omega},\mathbf{p})\big/\alpha_{i}(\psi_{i-1}(\mathbf{x}_{0},0,\boldsymbol{\omega},\mathbf{p})),\;0\Big\}$ (38)
@@
are enforced by applying Softplus activations to the corresponding output channels of $\mathfrak{N}$, ensuring that each $p_{i}^{\mathrm{inip}}$ satisfies its required lower bound. InitNet is only used at time $t=0$ to provide a set of HOCBFs and the initial multipliers for corresponding class $\kappa$ functions, which are then used to train $\bm{\theta}_{b}$ to obtain $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\mathrm{inip}})$.
@@
RefNet Structure. At each discrete time step, we employ a neural network (with or without memory), referred to as RefNet, parameterized by $(\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{q})$—to generate the reference terms $\mathbf{F}(\mathbf{x}_{0:t},\boldsymbol{\theta}_{f}),\mathbf{Q}(\mathbf{x}_{0:t},\boldsymbol{\theta}_{q})$ for controller $\pi(\mathbf{x}_{0:t},\bm{\theta})$ in (3) or $\mathbf{F}(\mathbf{x}(t),\boldsymbol{\theta}_{f}),\mathbf{Q}(\mathbf{x}(t),\boldsymbol{\theta}_{q})$ for controller $\pi(\mathbf{x}(t),\bm{\theta})$ in (2)—which serve as reference inputs to the BarrierNet module.
@@
BarrierNet Structure. In this paper, BarrierNet is implemented as a neural network that takes the system state $\mathbf{x}(t)$ as input and outputs the multipliers $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\mathrm{inip}})$ and the optimal control input $\mathbf{u}^{*}(t)$, where the network is parameterized by $\boldsymbol{\theta}_{p}$. The hyperparameters of the time-varying HOCBFs are obtained from the InitNet $\mathfrak{N}(\mathbf{x}_{0},\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{\mathrm{inip}})$, which produces the vectors $\boldsymbol{\Omega}$ and $\mathbf{P}_{\mathrm{init}}$ that encode all $\gamma$-function hyperparameters and the initial multipliers.
@@
Given the outputs of RefNet and InitNet, the BarrierNet layer solves the differentiable QP in (10) at each time step to produce the multipliers $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\mathrm{inip}})$ and the control input $\mathbf{u}^{*}(t)$. The control input $\mathbf{u}^{*}(t)$ is then applied sequentially to the system (1), allowing BarrierNet to roll out the closed-loop trajectory. Based on the resulting trajectory $\mathbf{x}_{0:t_{T}}$ and the corresponding multipliers $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\mathrm{inip}})$, the overall robustness $\rho(\tilde{\varphi},\mathbf{x}_{0:t_{T}},0)$ of the augmented STL specification, together with the associated cost $J(\mathbf{u}_{0:t_{T}-1})$, is evaluated. This trajectory-level evaluation is then used to update the controller parameters $\boldsymbol{\theta}$ during reinforcement training. Overall, the neural network controller $\pi(\cdot,\boldsymbol{\theta})$, where $\boldsymbol{\theta}=(\boldsymbol{\theta}_{q},\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{\text{inip}},\boldsymbol{\theta}_{p}),$ consists of (i) InitNet for initial hyperparameter generation, (ii) RefNet for predicting $\mathbf{F}$ and $\mathbf{Q}$, and (iii) the BarrierNet dQP layer that enforces the HOCBF constraints. The full architecture is illustrated in Fig. 1.
@@
Training Neural Network Controller. Following the procedure in [23], we randomly sample $V$ initial conditions $\mathbf{x}_{0}^{v}$, $v=1,\ldots,V$. For each initial state, we roll out the closed-loop system (1) under the controller $\pi$ up to the horizon $T$, thereby generating $V$ state and control trajectories. For each trajectory, we compute the unified STL robustness and its associated cost $J$, and use their empirical mean to approximate the expectation. Thus, the optimization problem in (8) can be written as
@@
where the superscript $v$ denotes the $v$-th sample. By substituting the system dynamics into the objective, the problem becomes an unconstrained optimization over the parameters of the controller. Since the QP in (10) is differentiable with respect to its inputs and parameters via the OptNet technique [4], we backpropagate the gradient of the objective in (39) through the QP layer to all components of $\bm{\theta}$. The gradients of the STL robustness are computed analytically and automatically using a modified implementation of STLCG [14] adapted to the exponential robustness in [21]. We then update the parameters using these gradients. At each gradient step, we resample $V$ initial conditions from the set $\mathcal{X}_{0}$ to improve exploration of the initial-state distribution, and we employ the stochastic optimizer Adam [13] to perform the training. We summarize our solution to Problem 1 in Algorithm 1.
@@ 2 Set up constraints on $\bm{\Omega}$ using (22), (24), (27);
3 Set up constraints on $\mathbf{P}_{\mathrm{inip}}$ using (9);
@@ V-A Case Study I: Linear Dynamical System — Double Integrator
where $\mathbf{x}=[x\ y\ v_{x}\ v_{y}]^{\top}$ and $\mathbf{u}=[u_{1}\ u_{2}]^{\top}$, with $x\ y]^{\top}$ denoting the 2D position, $v_{x}\ v_{y}]^{\top}$ the velocity, and $u_{1}\ u_{2}]^{\top}$ the acceleration of the robot. During the initialization of the neural network controller, control input bounds are enforced to ensure a proper warm start. However, in the subsequent optimization stage, the inputs are treated as unconstrained, while the STL specification still accounts for the actual input limits in (34), and the cost function in (39) employs an $\ell_{2}$ penalty with a weight of $0.003$ to discourage large accelerations.
@@ V-A1 Neural Network Controller without Memory
Comparison setup. We construct the HOCBFs (19), (20) and train the controller proposed in this paper under two different configurations. (i) In the first configuration, the multipliers $p_{i,j}$ are time-varying, while the feasibility subformulas (32), (33) are not incorporated; we refer to this controller as BN-VarP. (ii) In the second configuration, the multipliers $p_{i,j}$ are time-varying and the feasibility subformulas (32), (33) are incorporated; we refer to this controller as FeasiBN-VarP. The position state [xy][x\ y] serves as the input to the trained multiplier $p_{i,j}([x\ y],\boldsymbol{\theta}_{p},\mathbf{P}_{\text{init}})$.
For comparison, we include our previous work [24], where the multipliers $p_{i,j}$ remain constant within each rollout; we call this controller BN-FixedP. Another benchmark is [23], in which a neural network controller without BarrierNet—i.e., a standard Fully Connected Neural Network (FCNet)—is trained to satisfy an STL task. This is equivalent to directly using the reference control $\mathbf{F}$. We refer to this controller as FCNet. To ensure a fair comparison, we assume that the system dynamics are known for [23], and we use the same objective function, optimizer, and neural network architectures; that is, BN-FixedP and FCNet share the same architecture as $\mathbf{F}(\mathbf{x},\bm{\theta}_{f})$.
The training curves for all learning-based methods are shown in Fig. 3. In addition, we apply the method in [16] (extended to the HOCBF setting) without learning: we construct HOCBFs with fixed hyperparameters and solve the QP (10) with $\mathbf{F}=\mathbf{0}$. The hyperparameters are randomly chosen but satisfy all constraints (22), (24), and (27). We refer to this baseline as HOCBF. The resulting average objective values and robustness (evaluated only for the original task $\varphi$ and the control-bound specification $\varphi^{\mathcal{U}}$) over 10 random initial conditions are shown in Fig. 3, where the HOCBF baseline is illustrated by pink dashed lines. Representative trajectories obtained by the five approaches under 10 random initial conditions are shown in Fig. 2(b). The inputs over time for the five methods are shown in Fig. 4. We select the second task, $F_{[2,5]}Reg_{2}$ ($j=2$), and the fourth task, $G_{[0,5]}(\neg Obs_{4})$ ($j=4$), to illustrate the corresponding $p_{i,j}$ values ($i=1,2$) for the five methods in Fig. 5.
@@
Figure 3: Learning curves for BN-FixedP, BN-VarP, FeasiBN-VarP, and FCNet. Dashed lines show the results of directly using HOCBFs. (a) mean robustness $\rho(\varphi\wedge\varphi^{\mathcal{U}},\mathbf{x},0)$ during training. (b) mean objective values during training.
@@
Analysis and Discussion. From Fig. 3(a), we observe that when using BarrierNet (BN-FixedP, BN-VarP, and FeasiBN-VarP), all three methods achieve high robustness after sufficient training iterations. Specifically, BN-FixedP reaches a maximum robustness of approximately 0.45 during 500 iterations, while BN-VarP and FeasiBN-VarP reach about 0.6 and 0.7, respectively. Moreover, the robustness obtained by each BarrierNet-based method is higher than that of FCNet at nearly every iteration, demonstrating the correctness of Theorem 5. Occasionally, FCNet achieves higher robustness than BN-FixedP. This occurs because fixed multipliers limit the performance of BarrierNet, especially when the STL specification involves control bounds. In contrast, the use of time-varying multipliers in BN-VarP and FeasiBN-VarP significantly improves robustness, confirming that time-varying multipliers enhance the satisfaction of STL specifications, as claimed in Rem. 1. We further note that BN-VarP produces negative robustness at certain iterations, and further inspection reveals that the controller fails to satisfy the control-bound–related specification at those times. In comparison, FeasiBN-VarP maintains positive robustness throughout training, showing that the feasibility-aware formulation effectively guides the QP toward satisfying all constraints.
@@
In both Fig. 4 and Fig. 5, we plot the input and multiplier profiles from a single representative trajectory chosen from the 10 trajectories. From Fig. 4, we observe that after training, both the BarrierNet-based methods and FCNet produce control inputs that satisfy the input bounds. The input profiles of BN-FixedP closely resemble those of the HOCBF baseline after training, as Fig. 5 shows that the learned multipliers of BN-FixedP are constant over time and close to the fixed HOCBF multipliers after training. These results indicate that fixed multipliers substantially limit the ability of BarrierNet to improve performance. In contrast, BN-VarP and FeasiBN-VarP adjust the inputs earlier and more aggressively to satisfy the specifications, as shown in Fig. 4(b). This behavior is consistent with Fig. 5, where the time-varying multipliers adapt across time steps according to the control needs. Such adaptability enables these methods to achieve higher robustness.